OpenStudy (anonymous):

About an integer A, we know that 1. It has a remainder of 5 when being divided by 11, 2. A remainder of 4 when divided by 9 3. A remainder of 5 when divided by 7 4. A remainder of 4 when divided by 5 What is the remainder when A is divided by 3?

OpenStudy (amoodarya):

from 2 we know a=9q+4 so when divided by 3 we have a=3(3q)+4=3(3q)+3+1=3(3q+1)+1=3q'+1 so theremainder when A is divided by 3 is 1

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

How did I not think of that?